# Reasoning Language Models: A Blueprint

### v3

@@ Abstract
Reasoning language models (RLMs), also known as Large Reasoning Models (LRMs), such as OpenAI’s o1 and o3, DeepSeek-V3, and Alibaba’s QwQ, have redefined AI’s problem-solving capabilities by extending large language models (LLMs) with advanced reasoning mechanisms. Yet, their high costs, proprietary nature, and complex architectures—uniquely combining Reinforcement Learning (RL), search heuristics, and LLMs—present accessibility and scalability challenges. To address these, we propose a comprehensive blueprint that organizes RLM components into a modular framework, based on a survey and analysis of all RLM works. This blueprint incorporates diverse reasoning structures (chains, trees, graphs, and nested forms), reasoning strategies (e.g., Monte Carlo Tree Search, Beam Search), RL concepts (policy, value models and others), supervision schemes (Outcome-Based and Process-Based Supervision), and other related concepts (e.g., Test-Time Compute, Retrieval-Augmented Generation, agent tools). We also provide detailed mathematical formulations and algorithmic specifications to simplify RLM implementation. By showing how schemes like LLaMA-Berry, QwQ, Journey Learning, and Graph of Thoughts fit as special cases, we demonstrate the blueprint’s versatility and unifying potential. To illustrate its utility, we introduce x1, a modular implementation for rapid RLM prototyping and experimentation. Using x1 and a literature review, we provide key insights, such as multi-phase training for policy and value models, and the importance of familiar training distributions. Finally, we discuss scalable RLM cloud deployments and we outline how RLMs can integrate with a broader LLM ecosystem. Our work demystifies RLM construction, democratizes advanced reasoning capabilities, and fosters innovation, aiming to mitigate the gap between “rich AI” and “poor AI” by lowering barriers to RLM design and experimentation.
@@ 1 Introduction
Reasoning Language Models (RLMs), such as OpenAI’s o1 [116], o3 [76], and Alibaba’s QwQ [148], also referred to as Large Reasoning Models (LRMs) [^1] We use the term “Reasoning Language Model” instead of “Large Reasoning Model” because the latter implies that such models are always large. This does not necessarily have to be the case – as a matter of fact, smaller RLM can outperform larger LLMs [56]., represent a transformative breakthrough in AI, on par with the advent of ChatGPT [114]. These advanced systems have fundamentally redefined AI’s problem-solving capabilities, enabling nuanced reasoning, improved contextual understanding, and robust decision-making across a wide array of domains, reshaping science [45], industries [21], governance [52], and numerous other aspects of human life [80, 46, 75, 144, 143]. By extending the capabilities of standard large language models (LLMs) with sophisticated reasoning mechanisms, RLMs have emerged as the new cornerstone of cutting-edge AI, bringing us closer to AGI.
However, the high cost and proprietary nature of state-of-the-art RLMs, such as those developed by OpenAI, risk exacerbating the divide between “rich AI” and “poor AI”, raising significant concerns about accessibility and equity. Even the publicly available QwQ only comes with its model weights, and Alibaba does not disclose details about their training or data generation methodologies. Businesses and individuals unable to afford these advanced systems face a growing disadvantage, threatening to stifle innovation and reinforce systemic inequities. As RLMs become integral to critical applications, from healthcare to science, management, and beyond, it is imperative to address these disparities and ensure that the benefits of advanced reasoning capabilities are broadly accessible.
@@ 4.2.3 Reasoning Strategy
  • Beam Search [137] A breadth-limited search that keeps a fixed number of top-ranked continuations at each step. While commonly used for decoding token sequences, beam search can also apply to reasoning steps.
@@ 4.3.1 Structure Operators
Structure operators transform the reasoning structure by taking it as input and producing a modified version, typically through addition or refinement of reasoning steps. For instance, they may add new children to a specific node, facilitating the exploration of alternative reasoning paths.
@@
  • Refine The Refine operator enhances a given individual reasoning step. For example, it could address ambiguities, correct errors, and optimize inefficiencies, resulting in a more robust version of the step [99]. It could also integrate suggestions from self-critique [128] (evaluates steps to identify weaknesses and suggest targeted improvements), summarization [186] (condenses key elements into concise representations to streamline the reasoning structure), or rephrasing [43] (reformulates steps to improve clarity and coherence while preserving their logical integrity).
@@ 4.3.3 Update Operators
The Update operator enhances specific parts of the reasoning structure without altering the structure itself. A common example is the backpropagation phase in MCTS, where evaluation scores are propagated and updated along existing reasoning steps to inform future decisions. Another form of update involves refining the content of individual nodes or subsets of nodes, replacing their original versions with improved iterations, such as the “enhance” thought transformation in Graph of Thoughts [9].
@@ 4.4 Models
Models are further categorized and discussed in detail in Appendix B; we discuss the variants of the value model (Q Value model, V Value model), we compare Process Reward and Outcome Reward models, and we formally identify a new variant of models, the Outcome-Driven Process Reward Model.
@@ 4.4.2 Training Data Scope
OBS vs. PBS By varying the training data scope, developers can strike a balance between ease of data collection and the depth of reasoning insights provided to the model, with dense supervision generally offering improved performance at the cost of increased data complexity. We detail these, and additional aspects of ORMs and PRMs in Pipelines for different training phases and paradigms can be found in Appendix B, Appendix C.2, Appendix C.3, and in Algorithms 27.
@@ 4.5 Pipelines
Structure Step Strategy Gen. Ref. Agg. Pr. Res. Sel. BT Bp. Inter. Final. PM VM Inf. Tr. DG
@@ 5 Expressing Existing Schemes
We now showcase the expressivity of our blueprint, by illustrating how it can be used to model a broad scope of existing RLMs and other related works. We summarize the outcomes of the analysis in Table I. We start with typical and most prevalent Explicit RLM architectures based on MCTS and policy and/or value models, where a single reasoning step is an individual logical argument (Section 5.1). We also discuss there schemes that generalize this typical design, by harnessing nesting or Linearization Structure operators. Finally, we study Implicit RLMs (Section 5.2) and various structured prompting schemes such as Cot or ToT (Section 5.3), showing that they also fit our blueprint.
@@ 5.1 Explicit RLMs
Journey Learning [119] exhibits two main differences to typical MCTS-based RLMs. First, it harnesses the Linearization Structure operator, in which the tree reasoning structure is transformed into a chain, by extracting several selected reasoning chains from it and combining them together into an individual long chain. This way, the scheme attempts to harness insights from different tree branches. By maintaining a chain-based structure, Journey Learning preserves the simplicity of linear reasoning while embedding the capacity for self-correction and exploration of multiple hypotheses. Additionally, Journey Learning introduces a pipeline for the internalization of such long reasoning chains into its weights. This enables the final model to generate such long reasoning chains, possibly containing different reasoning branches, directly from its weights, making it an implicit RLM.
@@ 5.2 Implicit RLMs
Qwens’s QwQ [148] embodies a fully implicit reasoning model, characterized by an implicit reasoning structure that is generated autoregressively directly by the model weights. The reasoning strategy in QwQ – as indicated by the model output – harnesses next-step generation, backtracking, summarization, and critique generation to derive the final solution. At each step, the model implicitly generates a new node within the chain by employing one of these four implicit generate operators, presumably implemented using special tokens.
@@ 5.3 Structured Prompting Schemes
CoT [160] utilizes an implicit reasoning structure consisting of a chain of reasoning steps. The reasoning strategy employed in CoT is oriented towards constructing a single coherent chain of reasoning, culminating in a solitary solution, thus only needing the generation operator. CoT serves as the foundational framework for a range of advanced reasoning strategies, including prompting methodologies such as Self-Consistency and Self-Refinement, among others.
@@ 6.3 Part 3: Determine the Training Details
In this phase, you need to outline the specifics of training for the models that will implement operators. For an MCTS-based design, consider the typical approach of using the policy model to implement Generate (Expand) and the value model for Simulate. If necessary, you might also train a separate model to calculate the reward at individual nodes, enhancing the precision of the reward signals.
@@
Collect training data for both the policy and value models. For the policy model, consider generating data automatically with our pipeline or using a scheme such as CoT prompting, and include a special end-of-step token to ensure clean segmentation. For the value model, generate data through MCTS full simulations, which provide rich, structured information about reasoning paths and outcomes.
@@ 7.2 Operators
The framework implements two different Evaluate Operators. First, the Reasoning Path Evaluation operator predicts the discounted expected future reward for a chain extending from the root to a specific node. This prediction is derived from the q-value model, offering a quantitative measure of the path’s quality. Second, when the ground truth is available, the Ground Truth-Based Reward operator directly evaluates leaf nodes for correctness, assigning fixed rewards to verified solutions. These rewards are incorporated into the q-values of upstream nodes, ensuring that the reasoning process is informed by both model predictions and objective validation.
@@ 7.3 Models & Training Paradigms
Both the value and the policy model in x1 are fine-tuned versions of an LLM [^6] We currently use Llama-3.1-8B-Instruct as base model., without reliance on prompting, which is used in several other RLM architectures [56, 178]. This design decision aims to maximize the quality of results. We now outline briefly selected key aspects of how we train these models, full details can be found in Appendix B, C, and D.
@@ 7.3.2 Training the Value Model
$Q_{\pi}(s_{t},a_{t})$ $=\mathbb{E}\left[\gamma^{T-t}r(s_{T},a_{T})\mid s_{t},a_{t}\right]$ (1)
$\approx\frac{1}{N}\sum_{i=1}^{N}\gamma^{T-t}r(s_{T}^{(i)},a_{T}^{(i)})$ (2)
where $\gamma$ is the discount factor, $T$ marks the last reasoning step $a_{T}$ that is added resulting the terminal state $s_{T+1}$ containing the complete reasoning structure and rewards are modeled sparse. The terminal state $s_{T+1}$ is defined as the state in which no additional reasoning steps can be added. It typically represents the state containing the final solution to the problem at hand. Accordingly, $r(s_{T},a_{T})$ is the terminal reward. We chose to model rewards sparse, where only the final reasoning step receives a non-zero reward, since for most reasoning tasks, only the final answer can be evaluated against the true solution. As a result, one can only obtain a reward signal when the last step is reached. We can approximate the q-value by sampling $N$ reasoning chains until the terminal state, as in 2, and averaging the terminal rewards discounted by the depth required.
@@ 7.5 Blueprint for Efficient Scaling
  • Data Storage In the cloud, object storage provides automatic bandwidth scalability that allows scale computations operating on the same data. To overcome latency and power constraints, data can also be placed in in-memory caches like Redis and hybrid solutions that combine disks with flash memory [180].
@@
  • Parallelism We apply parallelism at multiple blueprint levels, including the classic data, model, and pipeline parallelism. These can scaled horizontally across a larger number of virtual machines and containers. On the other hand, reasoning steps can benefit from elastic scaling, like in distributed MCTS and Beam Search, where each path can be explored in parallel. There, containers can be allocated on the fly to support new paths and deallocated as soon as the parallelism scale of the computation decreases.
@@ 8 Example Insights for Effective RLMs
Use Two Phases for Training Adopting a two-phase training strategy—splitting SFT and RL—has proven effective in several contexts. This phased approach allows the model to first learn a solid foundation of reasoning patterns in phase one, followed by fine-tuning under more complex, adaptive conditions in phase two. For instance, research on Process Reinforcement through Implicit Rewards demonstrates that models trained with a dedicated SFT phase can maintain performance on standard benchmarks while achieving improved reasoning capabilities during RL. This separation also helps mitigate instability and ensures that each phase targets specific learning objectives, ultimately leading to more robust RLMs.
@@ 9.1 Mathematical Reasoning
GSM8K [36] consists of a train set (7,473 samples) and a test set (1,319 samples) of high-quality grade school-level mathematical word problems. Early breakthroughs in mathematical problem-solving by language models were achieved by training on the training subset of this benchmark.
@@ 9.2 Logical Reasoning
Other benchmarks include ProofWriter [145], FOLIO [58], WANLI [89], CLUTRR [136], Adversarial NLI [112], AbductionRules [170], and Adversarial ARCT [113].
@@ 9.6 Reasoning Utilities
Benchmarking capabilities of RLMs related to reasoning utilizies involve testing the capabilities of an RLM in how it acts as an agent. This includes benchmarks such as GAIA [106], WebArena [185], Mind2Web [42], WebShop [168], ALFWorld [132], AgentBench [91], AgentGym [162], and AgentBoard [22]. Another line of related benchmarks tests the RAG capabilites [26, 164, 98, 47].
@@ 10.2 Explicit Reasoning Models
The following works explore techniques that extend beyond basic mechanisms applied during pre-training or inference. These methods involve additional computation to iteratively refine reasoning paths, often increasing computational demands during training and/or inference.
Dong et al. [44] provide a taxonomy and survey of inference-time self-improvement methods, including independent, context-aware, and model-aided approaches. Guan et al. [55] propose verifier engineering, a post-training paradigm for foundation models involving three stages: Search, Verify, and Feedback, to enhance model outputs with scalable supervision signals. Zeng et al. [175] provide a comprehensive roadmap for reproducing OpenAI’s o1 reasoning model from a reinforcement learning perspective. Although the work thoroughly examines all core components: policy initialization, reward design, search, and learning, no implementation is provided. Various specific implementations of RLMs exist, we provide a summary in Table I. There are also other works related to Explicit RLMs, considering both coarse reasoning steps [163, 157] and fine reasoning steps [41, 163, 157].

### v4

@@ Abstract
Reasoning language models (RLMs), also known as Large Reasoning Models (LRMs), such as OpenAI’s o1 and o3, DeepSeek-R1, and Alibaba’s QwQ, have redefined AI’s problem-solving capabilities by extending large language models (LLMs) with advanced reasoning mechanisms. Yet, their high costs, proprietary nature, and complex architectures—uniquely combining reinforcement learning (RL), search heuristics, and LLMs—present accessibility and scalability challenges. To address these, we propose a comprehensive blueprint that organizes RLM components into a modular framework, based on a survey and analysis of all RLM works. This blueprint incorporates diverse reasoning structures (chains, trees, graphs, and nested forms), reasoning strategies (e.g., Monte Carlo Tree Search, Beam Search), RL concepts (policy, value models and others), supervision schemes (Outcome-Based and Process-Based Supervision), and other related concepts (e.g., Test-Time Compute, Retrieval-Augmented Generation, agent tools). We also provide detailed mathematical formulations and algorithmic specifications to simplify RLM implementation. By showing how schemes like LLaMA-Berry, QwQ, Journey Learning, and Graph of Thoughts fit as special cases, we demonstrate the blueprint’s versatility and unifying potential. To illustrate its utility, we introduce x1, a modular implementation for rapid RLM prototyping and experimentation. Using x1 and a literature review, we provide key insights, such as multi-phase training for policy and value models, and the importance of familiar training distributions. Finally, we discuss scalable RLM cloud deployments and we outline how RLMs can integrate with a broader LLM ecosystem. Our work demystifies RLM construction, democratizes advanced reasoning capabilities, and fosters innovation, aiming to mitigate the gap between “rich AI” and “poor AI” by lowering barriers to RLM design and experimentation.
@@ 1 Introduction
Reasoning Language Models (RLMs), such as OpenAI’s o1 [120], o3 [79], DeepSeek-R1 [92] and Alibaba’s QwQ [152], also referred to as Large Reasoning Models (LRMs) [^1] We use the term “Reasoning Language Model” instead of “Large Reasoning Model” because the latter implies that such models are always large. This does not necessarily have to be the case – as a matter of fact, smaller RLM can outperform larger LLMs [57]. , represent a transformative breakthrough in AI, on par with the advent of ChatGPT [118]. These advanced systems have fundamentally redefined AI’s problem-solving capabilities, enabling nuanced reasoning, improved contextual understanding, and robust decision-making across a wide array of domains, reshaping science [47], industries [23], governance [53], and numerous other aspects of human life [83, 48, 78, 148, 147]. By extending the capabilities of standard large language models (LLMs) with sophisticated reasoning mechanisms, RLMs have emerged as the new cornerstone of cutting-edge AI, bringing us closer to AGI.
However, the high cost and proprietary nature of state-of-the-art RLMs, such as those developed by OpenAI, risk exacerbating the divide between “rich AI” and “poor AI”, raising significant concerns about accessibility and equity. Even the publicly available QwQ only comes with its model weights, and Alibaba does not disclose details about their training or data generation methodologies. Businesses and individuals unable to afford these advanced systems face a growing disadvantage, threatening to stifle innovation and reinforce systemic inequities. As RLMs become integral to critical applications, from healthcare to science, manage-
ment, and beyond, it is imperative to address these disparities and ensure that the benefits of advanced reasoning capabilities are broadly accessible.
@@ 4.2.3 Reasoning Strategy
  • Beam Search [141] A breadth-limited search that keeps a fixed number of top-ranked continuations at each step. While commonly used for decoding token sequences, beam search can also be applied to reasoning steps.
@@ 4.3.1 Structure Operators
Structure operators transform the reasoning structure by taking it as input and producing a modified version, typically through addition of reasoning steps. For instance, they may add new children to a specific node, facilitating the exploration of alternative reasoning paths.
@@ 4.3.3 Update Operators
The update operator enhances specific parts of the reasoning structure without altering the structure itself. A common example is the backpropagation phase in MCTS, where evaluation scores are propagated and updated along existing reasoning steps to inform future decisions. Another form of update involves refining the content of individual nodes or subsets of nodes, replacing their original versions with improved iterations. The refine operator could address ambiguities, correct errors, and optimize inefficiencies, resulting in a more robust version of the step [103]. It could also integrate suggestions from self-critique [132] (evaluates steps to identify weaknesses and suggest targeted improvements), summarization [191] (condenses key elements into concise representations to streamline the reasoning structure), or rephrasing [45] (reformulates steps to improve clarity and coherence while preserving their logical integrity). An example would be the “improve” thought transformation in Graph of Thoughts [11].
@@ 4.4 Models
Models are further categorized and discussed in detail in Appendix B; we discuss the variants of the value model (Q-Value Model, V-Value Model), we compare Process-Based and Outcome-Based Reward Models, and we formally identify a new variant of models, the Outcome-Driven Process-Based Reward Model.
@@ 4.4.2 Training Data Scope
OBS vs. PBS By varying the training data scope, developers can strike a balance between ease of data collection and the depth of reasoning insights provided to the model, with dense supervision generally offering improved performance at the cost of increased data complexity. We detail these, and additional aspects of ORMs and PRMs in regards to pipelines for different training phases and paradigms in Appendix B, Appendix C.2, Appendix C.3, and in Algorithms 27.
@@
Structure Step Strategy Gen. Agg. Pr. Res. Sel. BT Ref. Bp. Inter. Term. PM VM Inf. Tr. DG
@@ 5 Expressing Existing Schemes
We now showcase the expressivity of our blueprint, by illustrating how it can be used to model a broad scope of existing RLMs and other related works. We summarize the outcomes of the analysis in Table I. We start with typical and most prevalent Explicit RLM architectures based on MCTS and policy and/or value models, where a single reasoning step is an individual logical argument (Section 5.1). We also discuss there schemes that generalize this typical design, by harnessing nesting or restructure operators. Finally, we study Implicit RLMs (Section 5.2) and various structured prompting schemes such as CoT or ToT (Section 5.3), showing that they also fit our blueprint.
@@ 5.1 Explicit RLMs
Journey Learning [123] exhibits two main differences to typical MCTS-based RLMs. First, it harnesses a linearization restructure operator, in which the tree reasoning structure is transformed into a chain, by extracting several selected reasoning chains from it and combining them together into an individual long chain. This way, the scheme attempts to harness insights from different tree branches. By maintaining a chain-based structure, Journey Learning preserves the simplicity of linear reasoning while embedding the capacity for self-correction and exploration of multiple hypotheses. Additionally, Journey Learning introduces a pipeline for the internalization of such long reasoning chains into its weights. This enables the final model to generate such long reasoning chains, possibly containing different reasoning branches, directly from its weights, illustrating the path towards the construction of an implicit RLM.
@@ 5.2 Implicit RLMs
Qwens’s QwQ [152] embodies a fully implicit reasoning model, characterized by an implicit reasoning structure that is generated autoregressively directly by the model weights. The reasoning strategy in QwQ – as indicated by the model output – harnesses next-step generation, backtracking, summarization, and critique generation to derive the final solution. At each step, the model implicitly generates a new node within the chain by employing one of these four implicit operators, presumably implemented using special tokens.
@@ 5.3 Structured Prompting Schemes
CoT [165] utilizes an implicit reasoning structure consisting of a chain of reasoning steps. The reasoning strategy employed in CoT is oriented towards constructing a single coherent chain of reasoning, culminating in a solitary solution, thus only needing the generate operator. CoT serves as the foundational framework for a range of advanced reasoning strategies, including prompting methodologies such as Self-Consistency and Self-Refinement, among others.
@@ 6.3 Part 3: Determine the Training Details
In this phase, you need to outline the specifics of training for the models that will implement operators. For an MCTS-based design, consider the typical approach of using the policy model to implement generate (expand) and the value model for evaluation. If necessary, you might also train a separate model to calculate the reward at individual nodes, enhancing the precision of the reward signals.
@@
Collect training data for both the policy and value models. For the policy model, consider generating data automatically with our pipeline or using a scheme such as CoT prompting, and include a special end-of-step token to ensure clean segmentation. For the value model, generate data through full MCTS simulations, which provide rich, structured information about reasoning paths and outcomes.
@@ 7.2 Operators
The framework implements two different evaluate operators. First, the value estimate operator predicts the discounted expected future reward for a chain extending from the root to a specific node. This prediction is derived from the q-value model, offering a quantitative measure of the path’s quality. Second, when the ground truth is available, the terminality evaluation operator directly evaluates leaf nodes for correctness, assigning fixed rewards to verified solutions. These rewards are incorporated into the q-values of upstream nodes, ensuring that the reasoning process is informed by both model predictions and objective validation.
@@ 7.3 Models & Training Paradigms
Both the value and the policy model in x1 are fine-tuned versions of an LLM [^6] We currently use Llama-3.1-8B-Instruct as base model., without reliance on prompting, which is used in several other RLM architectures [57, 183]. This design decision aims to maximize the quality of results. We now outline briefly selected key aspects of how we train these models, full details can be found in Appendices B, C, and D.
@@ 7.3.2 Training the Value Model
$Q_{\pi}(s_{t},a_{t})$ $=\mathbb{E}\left[\gamma^{T-t}r(s_{T},a_{T})\mid s_{t},a_{t}\right]$
$\approx\frac{1}{N}\sum_{i=1}^{N}\gamma^{T-t}r(s_{T}^{(i)},a_{T}^{(i)}),$ (1)
where $\gamma$ is the discount factor, $T$ marks the last reasoning step $a_{T}$ that is added, resulting in the terminal state $s_{T+1}$ containing the complete reasoning structure and rewards are modeled sparse. The terminal state $s_{T+1}$ is defined as the state in which no additional reasoning steps can be added. It typically represents the state containing the final solution to the problem at hand. Accordingly, $r(s_{T},a_{T})$ is the terminal reward. We chose to model rewards as sparse, where only the final reasoning step receives a non-zero reward, since for most reasoning tasks, only the final answer can be evaluated against the true solution. As a result, one can only obtain a reward signal when the last step is reached. We can approximate the q-value by sampling $N$ reasoning chains until the terminal state, as in Eq. 1, and averaging the terminal rewards discounted by the depth required.
@@ 7.5 Blueprint for Efficient Scaling
  • Data Storage In the cloud, object storage provides automatic bandwidth scalability that allows to scale computations operating on the same data. To overcome latency and power constraints, data can also be placed in in-memory caches like Redis and hybrid solutions that combine disks with flash memory [185].
@@
  • Parallelism We apply parallelism at multiple blueprint levels, including the classic data, model, and pipeline parallelism. These can scale horizontally across a larger number of virtual machines and containers. On the other hand, reasoning steps can benefit from elastic scaling, like in distributed MCTS and Beam Search, where each path can be explored in parallel. There, containers can be allocated on the fly to support new paths and deallocated as soon as the parallelism scale of the computation decreases.
@@ 8 Example Insights for Effective RLMs
Use Two Phases for Training Adopting a two-phase training strategy—splitting SFT and RL—has proven effective in several contexts. This phased approach allows the model to first learn a solid foundation of reasoning patterns in phase one, followed by fine-tuning under more complex, adaptive conditions in phase two. For instance, research on process reinforcement through implicit rewards [41] demonstrates that models trained with a dedicated SFT phase can maintain performance on standard benchmarks while achieving improved reasoning capabilities during RL. This separation also helps mitigate instability and ensures that each phase targets specific learning objectives, ultimately leading to more robust RLMs.
@@ 9.1 Mathematical Reasoning
GSM8K [38] consists of a training set (7,473 samples) and a test set (1,319 samples) of high-quality grade school-level mathematical word problems. Early breakthroughs in mathematical problem-solving by language models were achieved by training on the training subset of this benchmark.
@@ 9.2 Logical Reasoning
Other benchmarks include ProofWriter [149], FOLIO [60], WANLI [93], CLUTRR [140], Adversarial NLI [116], AbductionRules [175], PARARULE-Plus [9], ReClor-plus [8], LogiQA-plus [8], LogiQAv2-plus [8] and Adversarial ARCT [117].
@@ 9.6 Reasoning Utilities
Benchmarking capabilities of RLMs related to reasoning utilizies involve testing the capabilities of an RLM in how it acts as an agent. This includes benchmarks such as GAIA [110], WebArena [190], Mind2Web [44], WebShop [173], ALFWorld [136], AgentBench [95], AgentGym [167], and AgentBoard [24]. Another line of related benchmarks tests the RAG capabilities [28, 169, 102, 49].
@@ 10.2 Explicit Reasoning Models
The following works explore techniques that extend beyond basic mechanisms applied during pre-training or inference. These methods involve iteratively refining reasoning paths, often increasing computational demands during training and/or inference.
Dong et al. [46] provide a taxonomy and survey of inference-time self-improvement methods, including independent, context-aware, and model-aided approaches. Guan et al. [56] propose verifier engineering, a post-training paradigm for foundation models involving three stages: Search, Verify, and Feedback, to enhance model outputs with scalable supervision signals. Zeng et al. [180] provide a comprehensive roadmap for reproducing OpenAI’s o1 reasoning model from a reinforcement learning perspective. Although the work thoroughly examines all core components: policy initialization, reward design, search, and learning, no implementation is provided. Various specific implementations of RLMs exist, we provide a summary in Table I. There are also other works related to Explicit RLMs, considering both coarse-grained [168, 162] and fine-grained [43, 168, 162] reasoning steps.